The SGS Co. had $193,000 in taxable income. Use the rates from Table 2.3.

What is the average tax rate? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

Average tax rate               %

What is the marginal tax rate? (Enter your answer as a percent rounded to the nearest whole number, e.g., 32.)
Answer #1

Tax=15%*50000+25000*25%+25000*34%+93000*39%=58520

Average Tax Rate=58520/193000=30.3212%

Marginal Tax Rate=39%